Two sides of a garden bed are to be lined with logs. Three logs are needed for each side but 1 meter is cut off the third lotto make sure that the log doesn’t extend beyond the garden. The width of the garden is 3 meters and it’s area is 42 meters squared. How long is a whole log? (Hint: let x be the length of a whole log)
Answer by Theo(13342)

the area of the garden is 42 square meters.
the width is 3 meters.
the length is 42 / 3 = 14 meters.

the 3 logs total measurements will be 1 meter more than 14 = 15 meters.

divide 15 meters by 3 and you get 5 meters per log.

you will lay 3 logs on each side along the length of the garden.

the total measurement of the 3 logs is 15 meters.

you cut off 1 meter from the third log on each side and the total measurement is now 14 meters which just fits.

to use x in this problem, you would do the following.

x is the length of one log.

there will be 3 logs on each side of the length whose total length will be 1 meter more than required.

therefore you get 3x = L + 1.

solve for L to get L = 3x - 1.

the area of the garden is equal to L * W, where L is the lenth and W is the width and A is the area.

since the area is 42 and the width is 3, then A = L * W becomes 42 = L * 3.

solve for L to get L = 42 / 3 = 14.

since L = 3x - 1, you get 3x - 1 = 14.

add 1 to both sides of this equation to get 3x = 15.

solve for x to get x = 5.

the length of a whole log is equal to x which is equal to 5.

that's your solution.